• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Aniol

    Log entry time 08:57:36 on December18,2007

    Entry number 221803

    keyword=run 4031,+4mm Pb scan

    During run 4031(+4 mm Pb scan)the high power heater turned off. We were hitting a frame inside the target, apparently. We turned off the beam and loop3 temperature recovered. The large peak at +0.03 in Y target is smaller at run 4030 (+3mm)

    FIGURE 1

    FIGURE 2